An asset has one expected cash flow of $1200, which is expected to occur in 6 months and a second expected cash flow of $7600, which is expected to occur in 15 months. All payments are equally risky and thus they have the same discount rate. You've decided that the discount rate is 9%. According to these numbers, how much is the asset worth to you today?  Round your answer to the NEAREST DOLLAR.
Answer #1

Let the cash flow in month n be denoted by CFn

Given,
CF6 = 1200
CF15 = 7600

Monthly Interest rate = r = 0.09/12

Present Value of a future cash flow = CFn/(1+r)n

Hence, present value of future cash flow = CF6/(1+r)6 + CF15/(1+r)15 = 1200/(1+0.09/12)6 + 7600/(1+0.09/12)15 = 7941.58 = $7942